OpenStudy (anonymous):

Angle KIJ=(3x+2) and angle HIJ=(15x-7) and angle KIH =85degrees. Find angles KIJ and HIJ. Help please!!

OpenStudy (tkhunny):

This cannot be solved as stated. Is there a triangle or drawing in here, somewhere?

OpenStudy (anonymous):

There are multiple choice answers A. Angle KIJ =17 and HIJ=68 B.angle KIJ = 68 and HIJ= 17 C.angle KIJ = 50 and HIJ= 80 D angle KIJ=80 and HIJ= 5

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

its more like a circle to me.. so (3x+2) + (15x-7) + 85 = 360 --> 18x = 280 --> x= 280/18 KIJ is 3x+2 = 3(280/18) + 2 = 48.67 HIJ is 15x-7 = 15(280/18) -7 = 226.33 226.33+48.67+85=360 degrees. correct me if im wrong

OpenStudy (anonymous):

@apihiah well there is multiple choice answers up there

OpenStudy (anonymous):

@apihliah

OpenStudy (anonymous):

didnt see that coming lol

OpenStudy (anonymous):

ok if im not mistaken.. KIJ and HIJ are inside KIH so (3x+2) + (15x-7) = 85 --> 18x = 90 --> x= 90/18 KIJ is 3x+2 = 3(90/18) + 2 = 17 HIJ is 15x-7 = 15(90/18) -7 = 68 so, its an A i suppose..correct me if im wrong

OpenStudy (anonymous):

Oh my goodness thank you sooo much!!! I was stuck on that problem for ages!! @apihliah

OpenStudy (anonymous):

no problem.. kinda feel good that i can help. later then